I am trying to verify the Identity of: cot x - csc x = cos x-1/ sin x can you please help?
Answer by Alan3354(21605)

I am trying to verify the Identity of: cot x - csc x = cos x-1/ sin x can you please help?
----------------------
cot x - csc x = cos x-1/ sin x
cot(x) = cos(x)/sin(x) and csc(x) = 1/sin(x)
(cos/sin) - 1/sin = (cos - 1)/sin